You spend $35 setting up a lemonade stand. It costs you $0.30 in supplies. To make each glass you charge $.75. How many glasses of lemonade do you have to sell to break even?
1
Expert's answer
2020-02-03T09:23:47-0500

Let "x" be a number of glasses you need to sell to break even.

It means that:

"(0.75-0.30)x=35," where "0.75-0.30=0.35" is a profit for one glass of lemonade

"0.35x=35"

"x=\frac{35}{0.35}=100"


Answer: 100 glasses.